### v1

@@ 6.6 Physical consistency of the uncovered Navier-Stokes solutions
Results in Sec. 6.5 show that, for a generic training setup, we are able to achieve small relative error for a variety of degrees of non-convexity. Because the two-dimensional turbulent flow case is chaotic, infinitesimal perturbations $\sim O(\epsilon)$ exponentially grow in time, therefore, the residual $\mathcal{R}({\bm{u}}(\bm{\Omega},t)+{\epsilon}(\bm{\Omega},t);{\bm{p}})\gg\mathcal{R}({\bm{u}}(\bm{\Omega},t);{\bm{p}})$ where ${\epsilon}$ is a perturbation parameter. In this section, we analyse the physical properties of the solutions of the Navier-Stokes equation (Eq. 22) uncovered by the PC-CNN, ${\bm{\eta_{\theta}}}$. First, we show snapshots of the time evolution of the two-dimensional turbulent flow in Figure 5. These confirm that the model is learning a physical solution in time, as shown in the error on a log-scale in panels (iv). The parameterisation of the bias is fixed in this case with $k_{\phi}=7,\mathcal{M}=0.5$.
@@
Second, we analyse the statistics of the solution. The mean kinetic energy of the flow at each timestep is directly affected by the introduction of the bias. In the case of our strictly positive bias ${\bm{\phi}}$ (Eq. (2)), the mean kinetic energy is increased at every point. Results in Figure 6(a) show the kinetic energy time series across the time-domain for: the true state ${\bm{u}}$; the biased data ${\bm{\zeta}}$; and the network’s predictions ${\bm{\eta_{\theta}}}$. The chaotic nature of the solution can be observed from the erratic (but not random) evolution of the energy. We observe that the network predictions accurately reconstruct the kinetic energy trace across the simulation domain.
@@ 7 Results on reconstruction from sparse information
${\bm{u}}({\bm{\Omega}_{L}},t)\triangleq{\left.\kern-1.2pt{\bm{u}}({\bm{\Omega}_{H}},t)\vphantom{\big{|}}\right|^{{\bm{\Omega}_{L}}}},$ (23)
where $|^{{\bm{\Omega}_{L}}}$ is the corestriction which extracts points only in the relevant domain.
@@ 7.1 Comparison with standard upsampling
$e=\sqrt{\frac{\sum_{t\in\mathcal{T}}\lVert{\bm{u}}(\bm{\Omega_{H}},t)-{\bm{f_{\theta}}}({\bm{u}}(\bm{\Omega_{L}},t))\rVert_{\bm{\Omega_{H}}}^{2}}{\sum_{t\in\mathcal{T}}\lVert{\bm{u}}(\bm{\Omega_{H}},t)\rVert_{\bm{\Omega_{H}}}^{2}}}.$ (24)
Upon discarding the transient, a solution ${\bm{u}}({\bm{\Omega}_{H}},t)\in\mathbb{R}^{70\times 70}$ is generated by time-integration over $12\times 10^{3}$ time-steps, with $\Delta t=1\times 10^{-3}$. We extract the low-resolution solution ${\bm{u}}({\bm{\Omega}_{L}},t)\in\mathbb{R}^{10\times 10}$. We extract $2048$ samples at random from the time-domain of the solution, each sample consisting of $\tau=2$ consecutive time-steps. The adam optimiser (Kingma and Ba, 2015) is employed for training with a learning rate of $3\times 10^{-4}$. We take $\lambda_{P}=10^{3}$ as the regularisation factor for the loss, as shown in Eq. (7), and train for a total of $10^{3}$ epochs, which is empirically determined to provide sufficient convergence, as per the results of a hyperparameter study (result not shown).
[…]
@@
Although the relative error provides a notion of predictive accuracy, it is crucial to assess the physical characteristics of the reconstructed field (Pope and Pope, 2000). The energy spectrum, which is characteristic of turbulent flows, represents a multi-scale phenomenon where energy content decreases with the wavenumber. From the energy spectrum of the network’s prediction, ${\bm{f_{\theta}}}({\bm{u}}(\bm{\Omega}_{L},t))$, we gain physical insight into the multi-scale nature of the solution. Results in Figure 7 show that the energy content of the low-resolution field diverges from that of the high-resolution field, which is a consequence of spectral aliasing. Network realisations ${\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))$ are capable of capturing finer scales of turbulence compared to both interpolation approaches, prior to diverging from the true spectrum as $\lvert{\bm{k}}\rvert=18$. These results show the efficacy of the method, recovering in excess of $99\%$ of the system’s total energy. The physics loss, $\mathcal{L}_{P}$, enables the network to act beyond simple interpolation. The network is capable of de-aliasing, thereby inferring unresolved physics. Parametric studies show similar results across a range of scale factors $\kappa$ (shown in C).
@@ 8 Conclusions
We propose a physics-constrained convolutional neural network (PC-CNN) to solve inverse problems that are relevant to spatiotemporal partial differential equations (PDEs). First, we introduce the physics-constrained convolutional neural network (PC-CNN), which provides the means to compute the physical residual, i.e. the areas of the field in which the physical laws are violated by the prediction provided by the convolutional neural network (CNN). We formulate an optimisation problem by leveraging prior knowledge of the underlying physical system to regularise the predictions from the PC-CNN. This is augmented by a simple time-windowing approach for the computation of temporal derivative. Second, we apply the PC-CNN to solve an inverse problem in which we are given data that is contaminated by a spatially varying systematic error (i.e., bias, or epistemic uncertainty). The task is to uncover from the biased data the true state, which is the solution of the PDE. We numerically test the method to remove biases from data generated by three partial differential equations of increasing complexity (linear and nonlinear diffusion-convection, and Navier-Stokes). The PC-CNN successfully infers the true state, and the bias as a by-product, for a large range magnitudes and degrees of non-convexity. This demonstrates that the method is accurate and robust. Third, we apply the PC-CNN to solve an inverse problem in which we are given sparse information on the solution of a PDE. The task is to reconstruct the solution in space with high-resolution but without using the full high-resolution data in the training. We reconstruct the spatiotemporal chaotic solution of the Navier-Stokes equations on a high-resolution grid from a sparse grid containing only 2% of the information. We demonstrate that the proposed PC-CNN provides accurate physical results, both qualitatively and quantitatively, as compared to traditional interpolation methods, such as bi-cubic interpolation. For both inverse tasks, we investigate the physical consistency of the inferred solutions for the two-dimensional turbulent flow. We find that the PC-CNN predictions provide correct physical properties of the underlying partial differential equation (Navier-Stokes), such as energy spectrum. This work opens opportunities for solving inverse problems with nonlinear partial differential equations. Future work will be focused on experimental data.
9 Acknowledgements
@@
D. Kelshaw. and L. Magri. acknowledge support from the UK Engineering and Physical Sciences Research Council and ERC Starting Grant PhyCo 949388. L.M. acknowledges support from the grant EU-PNRR YoungResearcher TWIN ERC-PI_0000005.Codes are available at: https://github.com/magrilab/pisr; https://github.com/magrilab/picr; https://github.com/magrilab/kolsol.
Appendix A Pseudo-spectral discretisation of Navier-Stokes residual
$\mathcal{R}(\bm{\hat{\bm{u}}}_{k};{\bm{p}})=\left(\tfrac{d}{dt}+\nu\lvert{\bm{k}}\rvert^{2}\right)\bm{\hat{\bm{u}}}_{k}-{\bm{\hat{f}}}_{k}+{\bm{k}}\frac{{\bm{k}}\cdot{\bm{\hat{f}}}_{k}}{\lvert{\bm{k}}\rvert^{2}}-{\hat{\bm{g}}}_{k},$ (25)
with ${\bm{\hat{f}}}_{k}=-\left(\widehat{{\bm{u}}\cdot\nabla{\bm{u}}}\right)_{k}$, where nonlinear terms are handled pseudospectrally, employing the $\nicefrac{{2}}{{3}}$ de-aliasing rule to avoid unphysical culmination of energy at the high frequencies Canuto et al. (1988). A solution is produced by time-integration of the dynamical system with the explicit forward-Euler scheme, choosing a time-step $\Delta t$ that satisfies the Courant-Friedrichs-Lewy (CFL) condition. Initial conditions are generated by producing a random field scaled by the wavenumber, which retains the spatial structures of varying lengthscale in the physical domain (Ruan and McLaughlin, 1998).
Appendix B Architectural Details
@@
B.1 Uncovering solutions from biased data
For the bias removal task, we use a UNet architecture to extract relevant features from the biased field. The architecture, as well as the general training loop can be seen in Figure 8.
@@ B.1 Uncovering solutions from biased data
B.2 Reconstruction from sparse information
For the reconstruction task, a three-layer convolutional neural network is used for the model. Figure 9 provides an overview of the architecture, as well as the general training loop.
@@ B.2 Reconstruction from sparse information
Appendix C Effect of scale factor on reconstruction results

### v3

@@ 6.6 Physical consistency of the uncovered Navier-Stokes solutions
Results in Sec. 6.5 show that, for a generic training setup, we are able to achieve small relative error for a variety of degrees of non-convexity. Because the two-dimensional turbulent flow case is chaotic, infinitesimal perturbations $\sim O(\epsilon)$ exponentially grow in time, therefore, the residual $\mathcal{R}({\bm{u}}(\bm{\Omega},t)+{\epsilon}(\bm{\Omega},t);{\bm{p}})\gg\mathcal{R}({\bm{u}}(\bm{\Omega},t);{\bm{p}})$ where ${\epsilon}$ is a perturbation parameter. In this section, we analyse the physical properties of the solutions of the Navier-Stokes equation (Eq. 6.5) uncovered by the PC-CNN, ${\bm{\eta_{\theta}}}$. First, we show snapshots of the time evolution of the two-dimensional turbulent flow in Figure 5. These confirm that the model is learning a physical solution in time, as shown in the error on a log-scale in panels (iv). The parameterisation of the bias is fixed in this case with $k_{\phi}=7,\mathcal{M}=0.5$.
@@
Second, we analyse the statistics of the solution. The mean kinetic energy of the flow at each timestep is directly affected by the introduction of the bias. In the case of our strictly positive bias ${\bm{\phi}}$ (Eq. (2.2)), the mean kinetic energy is increased at every point. Results in Figure 6(a) show the kinetic energy time series across the time-domain for: the true state ${\bm{u}}$; the biased data ${\bm{\zeta}}$; and the network’s predictions ${\bm{\eta_{\theta}}}$. The chaotic nature of the solution can be observed from the erratic (but not random) evolution of the energy. We observe that the network predictions accurately reconstruct the kinetic energy trace across the simulation domain.
@@ 7 Results on reconstruction from sparse information
${\bm{u}}({\bm{\Omega}_{L}},t)\triangleq{\left.\kern-1.2pt{\bm{u}}({\bm{\Omega}_{H}},t)\vphantom{\big{|}}\right|^{{\bm{\Omega}_{L}}}},$ (7.1)
where $|^{{\bm{\Omega}_{L}}}$ is the corestriction which extracts points only in the relevant domain. While the data is sparse in space, a minimal temporal resolution is required for the computation of the time derivatives.
@@ 7.1 Comparison with standard upsampling
$e=\sqrt{\frac{\sum_{t\in\mathcal{T}}\lVert{\bm{u}}(\bm{\Omega_{H}},t)-{\bm{f_{\theta}}}({\bm{u}}(\bm{\Omega_{L}},t))\rVert_{\bm{\Omega_{H}}}^{2}}{\sum_{t\in\mathcal{T}}\lVert{\bm{u}}(\bm{\Omega_{H}},t)\rVert_{\bm{\Omega_{H}}}^{2}}}.$ (7.2)
Upon discarding the transient, a solution ${\bm{u}}({\bm{\Omega}_{H}},t)\in\mathbb{R}^{70\times 70}$ is generated by time-integration over $12\times 10^{3}$ time-steps, with $\Delta t=1\times 10^{-3}$. We extract the low-resolution solution ${\bm{u}}({\bm{\Omega}_{L}},t)\in\mathbb{R}^{10\times 10}$. We extract $2048$ samples at random from the time-domain of the solution, each sample consisting of $\tau=2$ consecutive time-steps. The adam optimiser (Kingma & Ba, 2015) is employed for training with a learning rate of $3\times 10^{-4}$. We take $\lambda_{P}=10^{3}$ as the regularisation factor for the loss, as shown in Eq. (4.1), and train for a total of $10^{3}$ epochs, which is empirically determined to provide sufficient convergence, as per the results of a hyperparameter study (result not shown).
Figure 7: Reconstruction from sparse information: physics-constrained convolutional neural network (PC-CNN) compared with traditional interpolation methods to reconstruct a solution from a sparse grid ${\bm{\Omega}_{L}\in\mathbb{R}^{10\times 10}}$ (100 points) to a high-resolution grid ${\bm{\Omega}_{H}\in\mathbb{R}^{70\times 70}}$ (4900 points). Panel (i) shows the low-resolution input, ${\bm{u}}({\bm{\Omega}_{L}},t)$; (ii) bi-linear interpolation, $\textit{BL}({\bm{u}}({\bm{\Omega}_{L}},t))$; (iii) bi-cubic interpolation, $\textit{BC}({\bm{u}}({\bm{\Omega}_{L}},t))$; (iv) true high-resolution field, ${\bm{u}}({\bm{\Omega}_{H}},t)$; (v) model prediction of the high-resolution field, ${\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))$; and (vi) energy spectra for each of the predictions. Vertical lines $f^{n}$ denote the Nyquist frequencies for spectral grid, $f^{n}_{\bm{\hat{\Omega}_{k}}}$, and for the low-resolution grid, $f^{n}_{\bm{\Omega_{L}}}$.
@@
Although the relative error provides a notion of predictive accuracy, it is crucial to assess the physical characteristics of the reconstructed field (Pope & Pope, 2000). The energy spectrum, which is characteristic of turbulent flows, represents a multi-scale phenomenon where energy content decreases with the wavenumber. From the energy spectrum of the network’s prediction, ${\bm{f_{\theta}}}({\bm{u}}(\bm{\Omega}_{L},t))$, we gain physical insight into the multi-scale nature of the solution. Results in Figure 7 show that the energy content of the low-resolution field diverges from that of the high-resolution field, which is a consequence of spectral aliasing. Network realisations ${\bm{f_{\theta}}}({\bm{u}}({\bm{\Omega}_{L}},t))$ are capable of capturing finer scales of turbulence compared to both interpolation approaches, prior to diverging from the true spectrum as $\lvert{\bm{k}}\rvert=18$. These results show the efficacy of the method, recovering in excess of $99\%$ of the system’s total energy. The physics loss, $\mathcal{L}_{P}$, enables the network to act beyond simple interpolation. The network is capable of de-aliasing, thereby inferring unresolved physics. Parametric studies show similar results across a range of scale factors $\kappa$ (shown in C).
@@ 8 Conclusions
We propose a physics-constrained convolutional neural network (PC-CNN) to solve inverse problems that are relevant to spatiotemporal partial differential equations (PDEs). First, we introduce the physics-constrained convolutional neural network (PC-CNN), which provides the means to compute the physical residual, i.e. the areas of the field in which the physical laws are violated by the prediction provided by the convolutional neural network (CNN). We formulate an optimisation problem by leveraging prior knowledge of the underlying physical system to regularise the predictions from the PC-CNN. This is augmented by a simple time-windowing approach for the computation of temporal derivative. Second, we apply the PC-CNN to solve an inverse problem in which we are given data that is contaminated by a spatially varying systematic error (i.e., bias, or epistemic uncertainty). The task is to uncover from the biased data the true state, which is the solution of the PDE. We numerically test the method to remove biases from data generated by three partial differential equations of increasing complexity (linear and nonlinear diffusion-convection, and Navier-Stokes). The PC-CNN successfully infers the true state, and the bias as a by-product, for a large range magnitudes and degrees of non-convexity. This demonstrates that the method is accurate and robust. Third, we apply the PC-CNN to solve an inverse problem in which we are given sparse information on the solution of a PDE. The task is to reconstruct the solution in space with high-resolution but without using the full high-resolution data in the training. We reconstruct the spatiotemporal chaotic solution of the Navier-Stokes equations on a high-resolution grid from a sparse grid containing only $\lesssim 1\%$ of the information. We demonstrate that the proposed PC-CNN provides accurate physical results, both qualitatively and quantitatively, as compared to traditional interpolation methods, such as bi-cubic interpolation. For both inverse tasks, we investigate the physical consistency of the inferred solutions for the two-dimensional turbulent flow. We find that the PC-CNN predictions provide correct physical properties of the underlying partial differential equation (Navier-Stokes), such as energy spectrum. This work opens opportunities for solving inverse problems with nonlinear partial differential equations. Future work will be focused on experimental data.
Acknowledgments
The authors are grateful to G. Rigas for insightful discussions on turbulence.
Funding Statement
@@ Acknowledgments
D.Kelshaw. & L.Magri. acknowledge support from the UK Engineering & Physical Sciences Research Council, and ERC Starting Grant PhyCo 949388. L.M. acknowledges support from the grant EU-PNRR YoungResearcher TWIN ERC-PI_0000005.
Data Availability Statement
Codes and data are accessible via GitHub: https://github.com/MagriLab/pisr; https://github.com/MagriLab/picr; and https://github.com/MagriLab/kolsol.
Appendix A Architectural Details
@@
A.1 Uncovering solutions from biased data
For the bias removal task, we use a UNet architecture to extract relevant features from the biased field. The architecture, as well as the general training loop can be seen in Figure 8. This particular network configuration was found empirically. The input and output layers each feature two channels, corresponding to the two velocity components of the flow being studied. The network does not overfit, also thanks to the physics-based regularisation term in the loss function.
@@ A.1 Uncovering solutions from biased data
A.2 Reconstruction from sparse information
For the reconstruction task, a three-layer convolutional neural network is used for the model. Figure 9 provides an overview of the architecture, as well as the general training loop. The baseline network configuration is similar to that of Dong et al. (2014), with the additional constraints imposed through the physics-based loss function. The input and output layers each feature two channels, corresponding to the two velocity components of the flow being studied.
@@ A.2 Reconstruction from sparse information
Appendix B Pseudo-spectral discretisation of Navier-Stokes residual
$\mathcal{R}(\bm{\hat{\bm{u}}}_{k};{\bm{p}})=\left(\tfrac{d}{dt}+\nu\lvert{\bm{k}}\rvert^{2}\right)\bm{\hat{\bm{u}}}_{k}-{\bm{\hat{f}}}_{k}+{\bm{k}}\frac{{\bm{k}}\cdot{\bm{\hat{f}}}_{k}}{\lvert{\bm{k}}\rvert^{2}}-{\hat{\bm{g}}}_{k},$ (B.1)
with ${\bm{\hat{f}}}_{k}=-\left(\widehat{{\bm{u}}\cdot\nabla{\bm{u}}}\right)_{k}$, where nonlinear terms are handled pseudospectrally, employing the $\nicefrac{{2}}{{3}}$ de-aliasing rule to avoid unphysical culmination of energy at the high frequencies Canuto et al. (1988). A solution is produced by time-integration of the dynamical system with the explicit forward-Euler scheme, choosing a time-step $\Delta t$ that satisfies the Courant-Friedrichs-Lewy (CFL) condition. Initial conditions are generated by producing a random field scaled by the wavenumber, which retains the spatial structures of varying lengthscale in the physical domain (Ruan & McLaughlin, 1998).
Appendix C Parametric study on the scale factor